# Physical Processes Behind the Co-Evolution of Halos, Galaxies and Supermassive Black Holes in the IllustrisTNG Simulation

### v2

@@ 4.4 Phase-4: Merger dominated growth
The final stage of the evolution is a quenched phase. In this phase, the in-situ growth of SMBHs and galaxies ceases, while the ex-situ growth driven by galaxy-galaxy and SMBH-SMBH mergers continues. In the top left panel of Fig. 10, we show the distribution of $f_{\rm*,ex}^{[t_{\rm q},t_{z=0}]}$, the fraction of ex-situ growth of stellar mass between $t_{\rm q}$ and $z=0$, defined similarly to Eq. (20) but with a different time interval, for halos with different $M_{{\rm h},z=0}$. For comparison, in the other three panels, we show the distributions of the fraction of in-situ growth of stellar mass (bottom left), and the ex- and in-situ growth of SMBH mass (right column), respectively. For the ex-situ growth of stellar mass, a clear trend with halo mass is seen, with more massive halos having a larger fraction of ex-situ growth after quenching. This is expected as a consequence of the larger number of satellite galaxies in more massive halos (e.g. Yang et al., 2009, 2012; Wang et al., 2024c) carried in by halo-halo mergers. For halos with $M_{{\rm h},z=0}\geqslant 10^{12.6}\,{\rm M_{\odot}}$, there are cases where the ex-situ fraction exceeds $80\%$, indicating that the final properties of the central galaxies in these halos are shaped by their merger histories. These merger-dominated galaxies are expected to form in massive halos with late formation times, as shown by Chen et al. (2023) using a N-body simulation constrained by the density field that matches the observed structures found in SDSS (Wang et al., 2016). A plausible example is the Coma cluster, whose central subhalo appears at $z\approx 5$, a relatively late time in comparison with other massive halos, and then grows rapidly via a large number of major mergers. The BCG of Coma is thus likely to reside in the tail of the distribution of $f_{\rm*,ex}^{[t_{\rm q},t_{z=0}]}$. On the other hand, the fraction of in-situ growth of stellar mass since $t_{\rm q}$ is nearly independent of halo mass, and nearly always below $20\%$. Earlier works have suggested that significant in-situ growth of stellar mass can be induced by mergers, via the perturbation of angular momentum and the compression of gas by shocks (e.g. Capelo & Dotti, 2017; Lagos et al., 2018). Our result here suggests that, in TNG, the in-situ growth of stellar mass in galaxies that have already become “dry” (gas-poor) is insignificant.
@@
The distribution of in- and ex-situ fractions of $M_{\rm BH}$ growth is similar to those of $M_{*}$. Specifically, the ex-situ fraction of $M_{\rm BH}$ growth increases with $M_{{\rm h},z=0}$, and in extreme cases, can exceed $80\%$. The in-situ fraction of $M_{\rm BH}$ growth is always small, nearly zero, due to the strong kinetic-mode feedback that clears the gas surrounding the SMBH. The result here is consistent with Weinberger et al. (2018, see their §4.1 and Fig.7), where the growth of SMBHs in massive halos experiences a transition, from being dominated by in-situ accretion during the phase of thermal-mode feedback to being dominated by mergers during the phase of kinetic-mode feedback. As mergers tend to bring galaxies into a 1:1 relation in the $M_{\rm BH}$-$M_{*}$ plane, the $M_{\rm BH}$-$M_{*}$ scaling relation is expected to be modified during Phase-4. We will discuss this in §5.1.
@@
As the kinetic-mode AGN feedback in TNG is designed to sustain the quenching of massive galaxies via the self-regulated accretion (Weinberger et al., 2017), the sufficient condition (Eq. 21) of feedback-cooling balance for quenching should remain valid over the entire phase of the quenched evolution. The black contour in Fig. 9 shows a comparison of the cooling rate within the subhalo and the feedback energy rate at $z=0$. Both rates are shifted to lower values relative to those at $t_{\rm q}$, but remain around the balance line. Indeed, observations have suggested that the atmosphere of massive galaxies is in a state of marginally stable to form cold clouds, and the baryon cycling induced by self-regulating AGN feedback is capable of explaining the observed quasi-equilibrium state of massive galaxies (see, e.g. Donahue & Voit, 2022, for a review). A similar conclusion was reached by Zinger et al. (2020) using TNG, where gas entropy is used to quantify the effect of AGN feedback. Since high-entropy gas can have high temperature and thus high specific internal energy to be dispersed before forming stars, and have low cooling rate due to the high resistance to compression, the use of entropy as a proxy for the feedback effect is intuitive. Their results show that the entropy of the gas in the circumgalactic medium (CGM) of massive galaxies is significantly lifted by the kinetic-mode feedback out to a distance of several hundred kiloparsecs, so that quenching can be sustained. An interesting side effect of the kinetic-mode feedback has been demonstrated by Martín-Navarro et al. (2021), who found that the halo-scale wind driven by kinetic-mode feedback in TNG preferentially clears gas along the minor axes of the galaxies. This anisotropy of feedback reduces the effect of ram pressure stripping on the quenching of satellite galaxies along the minor axes of the central galaxies, and leads to an anisotropic quenching fraction of satellite galaxies that is consistent with the observational results from SDSS.
@@ 5.1 The build-up of scaling relations
Figure 11: $M_{\rm BH}$-$M_{*}$ (left panel) and $M_{\rm BH}$-$M_{\rm h}$ (right panel) relations at different time points, $t_{23}$ (blue contours), $t_{\rm q}$ (red contours), and $z=0$ (black dots), in the histories of the galaxies. In each panel, three dashed lines indicate three relations with different logarithmic slopes. Contours from inner to outer enclose 10% – 95% of the sample. See §5.1 for the details about the build-up of these relations.
As discussed in §4, the evolution of a SMBH (and its host galaxy/halo) involves a series of critical times that mark the transitions under specific conditions. In the $M_{\rm BH}$-$M_{*}$ (or $M_{\rm BH}$-$M_{\rm h}$) plane, these conditions form a series of edges that a SMBH has to pass before it can reach subsequent phases. SMBHs with different initial conditions start at different positions in the plane, “flow” in the plane following different paths, and cross the edges at different locations and times. Thus, the $M_{\rm BH}$-$M_{*}$ (or $M_{\rm BH}$-$M_{\rm h}$) scaling relations at a given epoch for a population of SMBHs in a given phase can be viewed as a time slice of the flow lines in the plane between the two edges that bracket the phase in question. If observations are able to obtain such slices of different redshifts, the locations of the edges can be inferred, thus providing constraints on the physical processes that drive the transitions in the evolution.
Fig. 11 shows the distributions of SMBHs in the $M_{\rm BH}$-$M_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ planes at three times: $t_{23}$, when the fast accretion of SMBH ends, $t_{\rm q}$ ($\equiv t_{34}$), when the galaxy is quenched, and $z=0$. As discussed in §4.2, the transition at $t_{23}$ happens when the thermal-mode feedback of AGN is able to unbind the gas of the galaxy. According to Eqs. (14) and (4.3.1), this corresponds to a condition of
@@
which predicts a power-law relation between $M_{\rm BH}$ and $M_{*}$ (or $M_{\rm h}$), with a power index $\gamma=5/3$ at the time when the star formation efficiency peaks, or a different index if $f_{\rm gas}$ and $f_{*}$ are mass-dependent. As shown by the blue contours in the left panel of Fig. 11, the $M_{\rm BH}$-$M_{*}$ relation at $t_{23}$ follows a power-law with a power index of about $5/3$. This implies a nearly constant $f_{\rm gas}f_{*}^{-5/3}$ at around $t_{\rm 23}$, and suggests that the gas and stellar masses are positively correlated. At $M_{*}\approx 10^{10}\,{\rm M_{\odot}}$ and $M_{\rm BH}\approx 10^{7}\,{\rm M_{\odot}}$, there is a population of SMBHs that deviate from the relation. As noted in §3, this is mainly due to the difficulty in determining $t_{23}$ accurately. On the other hand, as shown by the blue contours in the right panel, the $M_{\rm BH}$-$M_{\rm h}$ relation in logarithmic scales appears to be non-linear even for SMBHs with $M_{\rm BH}$ well above $10^{7}\,{\rm M_{\odot}}$. This reflects the non-linear dependence of $f_{\rm gas}$ on $M_{\rm h}$. At about $M_{\rm h}=M_{\rm h,c}=10^{12}\,{\rm M_{\odot}}$, $f_{\rm gas}$ reaches its peak, leading to a $M_{\rm BH}$-$M_{\rm h}$ relation that follows the power-law of $\gamma=5/3$. At lower or higher $M_{\rm h}$, the dependence of $f_{\rm gas}$ on $M_{h}$ changes, which results in bending in the $M_{\rm BH}$-$M_{\rm h}$ relation.
@@
As concluded in §4.3, the evolution of SMBHs between $t_{23}$ and $t_{\rm q}$ is self-regulated so as to follow $M_{\rm BH}\propto M_{*}^{5/3}$. Naively one would expect that the $M_{\rm BH}$-$M_{*}$ relation at $t_{\rm q}$ follows the same power law. However, as shown by the red contours in the left panel of Fig. 11, the $M_{\rm BH}$-$M_{*}$ relation at $t\sim t_{\rm q}$ appears to be much shallower, with $\gamma<1$. As discussed in §4.3.2, the trigger of the kinetic-mode feedback is controlled by the $M_{\rm BH}$-dependent threshold of the Eddington ratio, $\chi_{\rm thres}$ for SMBHs with $M_{\rm BH}\lesssim 7\times 10^{8}\,{\rm M_{\odot}}$. It can also be seen from Fig. 8 that the majority of galaxies are quenched soon (within $\approx 1\,{\rm Gyr}$) after the trigger of the kinetic-mode feedback for halos with $M_{\rm h}<10^{12.3}\,{\rm M_{\odot}}$. In TNG, $\chi_{\rm thres}$ is set so that the kinetic-mode feedback is triggered preferentially for high-mass SMBHs. Thus, the quenching of high-mass galaxies occurs earlier than that expected from a model with a constant (and small) $\chi_{\rm thres}$. The shallow $M_{\rm BH}$-$M_{*}$ relation at $M_{\rm*}\lesssim 10^{11}\,{\rm M_{\odot}}$ is a direct consequence of the choice of the $\chi_{\rm thres}$-$M_{\rm BH}$ relation. For higher masses ($M_{\rm BH}\gtrsim 7\times 10^{8}\,{\rm M_{\odot}}$), the threshold $\chi_{\rm thres}$ reaches the upper limit, 0.1, and becomes independent of $M_{\rm BH}$. Meanwhile, the conditions necessary for the quenching of galaxies in high-mass halos are delayed by frequent mergers (§4.3.3). Eventually, the preference for quenching for high-mass galaxies is weakened, which allows the SMBHs in them to grow further. The tip present in the red contours at $M_{*}\gtrsim 10^{11}\,{\rm M_{\odot}}$ in the left panel of Fig. 11 verifies this. The $M_{\rm BH}$-$M_{\rm h}$ relation at $t_{\rm q}$, as shown in the right panel, appears as shallow as $\gamma=0.5$. The difference in the index between $M_{\rm BH}$-$M_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ relations at $t_{\rm q}$ can be understood by an argument similar to that used for the bending $M_{\rm BH}$-$M_{\rm h}$ relation at $t_{\rm 23}$, but here both $f_{\rm gas}$ and $f_{*}$ decrease with increasing $M_{\rm h}$ at $M_{\rm h}\gtrsim 10^{12}\,{\rm M_{\odot}}$.
@@
The evolution of SMBHs (and their host galaxies) in the final phase after $t_{\rm q}$ is dominated by the ex-situ growth. As shown in §4.4, the fraction of ex-situ growth for both $M_{\rm BH}$ and $M_{*}$ depends significantly on the halo mass at $z=0$. This mass dependence implies that the growth of SMBHs and galaxies in low-mass halos is almost frozen after $t_{\rm q}$, while the growth in massive halos is not. The net effect is a “stretch” of the $M_{\rm BH}$-$M_{*}$ relation towards the high-mass end from an anchor point at the low-mass end, as shown by the black dots in the left panel of Fig. 11. As mergers tend to give a 1:1 relation between the involved masses, the $M_{\rm BH}$-$M_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ relations at the most massive ends steepen relative to those at $t_{\rm q}$. The history of the scaling relations is then erased by mergers, leaving behind little trace of the seeding and early growth of the SMBHs.
@@
The TNG subgrid model for AGN feedback is designed to reproduce the observed $M_{\rm BH}$-$M_{\rm*,bulge}$ relation of Kormendy & Ho (2013), which is a power-law form with an index of $\gamma\approx 1$. As shown by Weinberger et al. (2017, see their figure 5), the $M_{\rm BH}$-$M_{\rm*,bulge}$ relation of the TNG model roughly follows the $\gamma=1$ relation, and so does the $M_{\rm BH}$-$M_{*}$ relation. Since current observational results still have significant uncertainties in estimating the SMBH and bulge masses, it is still unclear whether or not the observed $M_{\rm BH}$-$M_{*}$ (or $M_{\rm*,bulge}$) relation can be extended to lower SMBH masses, and what the power index really is. In a more recent work, Greene et al. (2020) suggested a steeper relation, $M_{\rm BH}\sim M_{*}^{1.33}$ for early-type galaxies. Graham & Sahu (2023a, see their table 2) obtained $M_{\rm BH}\sim M_{*}^{1.69}$ (and $M_{\rm BH}\sim M_{\rm*,bulge}^{1.64}$) for pure elliptical galaxies. For the $M_{\rm BH}$-$M_{\rm h}$ relation, Voit et al. (2024) showed that the prediction from TNG is significantly flatter than observations. For the model to match the observations, they suggested modifications of the criteria for the triggering of the kinetic-mode feedback and/or the feedback efficiency. These more recent observations in the nearby Universe can thus put constraints on the subgrid physics to be implemented in the next generation of simulations, especially for the SMBH growth in the phase of quenched evolution.
@@
Since mergers in the quenched phase erased the imprints of early evolution, in particular for SMBHs in massive halos, a slice of the $M_{\rm BH}$-$M_{*}$ (or $M_{\rm BH}$-$M_{\rm h}$) flow lines in the nearby Universe may not be adequate to infer the full tracks of galaxies. This can be seen by comparing our TNG-based conclusions to those inferred by Graham & Sahu (2023a, see their figures 8 and 9) from the observed $M_{\rm BH}$-$M_{\rm*,bulge}$ relations for galaxies with different morphological types. At $M_{\rm*,bulge}\gtrsim 10^{11}\,{\rm M_{\odot}}$, their analysis suggests that mergers between elliptical galaxies (with bulge-to-total ratio $\approx 1$) tend to drive migrations in the $M_{\rm BH}$-$M_{\rm*,bulge}$ plane towards a 1:1 relation, consistent with our inferences from the TNG. However, our analysis suggests that, for galaxies with $M_{*}\lesssim 10^{10.5}\,{\rm M_{\odot}}$, the ex-situ growth in both $M_{*}$ and $M_{\rm BH}$ contributes little to the total mass, which is inconsistent with the merger-driven scenario for the formation of elliptical galaxies and the build-up of the $M_{\rm BH}$-$M_{\rm*,bulge}$ relation in the low-mass end. One possibility is that the dynamical hotness of these galaxies comes from the fast inflow of gas associated with the fast regime of halo assembly (e.g. Zhao et al., 2003), as suggested by Mo et al. (2024, see also ). The change of the power index of the $M_{\rm BH}$-$M_{*}$ relation at $t_{\rm q}$, as demonstrated in this paper, is then expected to occur if the kinetic-mode feedback drives the quenching of galaxies.
@@
On the other hand, observations directly targeting SMBHs at high $z$, where they are less massive and more active, are crucial to cover the early transitions/edges in the evolutionary history of the SMBH growth and to avoid the ambiguity in the extrapolation of the local scaling relations. Using the deep imaging and spectroscopy capabilities of JWST, Wang et al. (2024b) found a galaxy with a spectrum compatible with that produced by a Pop-III star cluster. As theoretical studies (e.g. Stacy et al., 2016; Wise et al., 2019; Klessen, 2019; Chon & Omukai, 2020; Latif et al., 2022a, b; Regan & Volonteri, 2024; Ventura et al., 2024) have suggested that massive Pop-III stars are promising SMBH seeds of intermediate masses ($\sim 10^{4}\,{\rm M_{\odot}}$), direct observations of Pop-III stars (clusters) will be crucial to distinguish among different seeding scenarios (see, e.g. Spinoso et al., 2023) and to probe the growth of SMBHs in Phase-1.

### v3

@@ 4.4 Phase-4: Merger dominated growth
The final stage of the evolution is a quenched phase. In this phase, the in-situ growth of SMBHs and galaxies ceases, while the ex-situ growth driven by galaxy-galaxy and SMBH-SMBH mergers continues. In the top left panel of Fig. 10, we show the distribution of $f_{\rm*,ex}^{[t_{\rm q},t_{z=0}]}$, the fraction of ex-situ growth of stellar mass between $t_{\rm q}$ and $z=0$, defined similarly to Eq. (20) but with a different time interval, for halos with different $M_{{\rm h},z=0}$. For comparison, in the other three panels, we show the distributions of the fraction of in-situ growth of stellar mass (bottom left), and the ex- and in-situ growth of SMBH mass (right column), respectively. For the ex-situ growth of stellar mass, a clear trend with halo mass is seen, with more massive halos having a larger fraction of ex-situ growth after quenching. This is expected as a consequence of the larger number of satellite galaxies in more massive halos (e.g. Yang et al., 2009, 2012; Wang et al., 2024c) carried in by halo-halo mergers. For halos with $M_{{\rm h},z=0}\geqslant 10^{12.6}\,{\rm M_{\odot}}$, there are cases where the ex-situ fraction exceeds $80\%$, indicating that the final properties of the central galaxies in these halos are shaped by their merger histories. These merger-dominated galaxies are expected to form in massive halos with late formation times, as shown by Chen et al. (2023) using an N-body simulation constrained by the density field that matches the observed structures found in SDSS (Wang et al., 2016). A plausible example is the Coma cluster, whose central subhalo appears at $z\approx 5$, a relatively late time in comparison with other massive halos, and then grows rapidly via a large number of major mergers. The BCG of Coma is thus likely to reside in the tail of the distribution of $f_{\rm*,ex}^{[t_{\rm q},t_{z=0}]}$. On the other hand, the fraction of in-situ growth of stellar mass since $t_{\rm q}$ is nearly independent of halo mass, and nearly always below $20\%$. Earlier works have suggested that significant in-situ growth of stellar mass can be induced by mergers, via the perturbation of angular momentum and the compression of gas by shocks (e.g. Capelo & Dotti, 2017; Lagos et al., 2018). Our result here suggests that, in TNG, the in-situ growth of stellar mass in galaxies that have already become “dry” (gas-poor) is insignificant.
@@
The distribution of in- and ex-situ fractions of $M_{\rm BH}$ growth is similar to those of $M_{*}$. Specifically, the ex-situ fraction of $M_{\rm BH}$ growth increases with $M_{{\rm h},z=0}$, and in extreme cases, can exceed $80\%$. The in-situ fraction of $M_{\rm BH}$ growth is always small, nearly zero, due to the strong kinetic-mode feedback that clears the gas surrounding the SMBH. The result here is consistent with Weinberger et al. (2018, see their §4.1 and Fig.7), where the growth of SMBHs in massive halos experiences a transition, from being dominated by in-situ accretion during the phase of thermal-mode feedback to being dominated by mergers during the phase of kinetic-mode feedback. As mergers tend to bring galaxies into a 1:1 relation in the $_{\rm BH}-MM_{*}$ plane, the $_{\rm BH}-MM_{*}$ scaling relation is expected to be modified during Phase-4. We will discuss this in §5.1.
@@
As the kinetic-mode AGN feedback in TNG is designed to sustain the quenching of massive galaxies via the self-regulated accretion (Weinberger et al., 2017), the sufficient condition (Eq. 21) of feedback-cooling balance for quenching should remain valid over the entire phase of the quenched evolution. The black contour in Fig. 9 shows a comparison of the cooling rate within the subhalo and the feedback energy rate at $z=0$. Both rates are shifted to lower values relative to those at $t_{\rm q}$, but remain around the balance line. Indeed, observations have suggested that the atmosphere of massive galaxies is in a state of marginally stable to form cold clouds, and the baryon cycling induced by self-regulating AGN feedback is capable of explaining the observed quasi-equilibrium state of massive galaxies (see, e.g. Donahue & Voit, 2022, for a review). A similar conclusion was reached by Zinger et al. (2020) using TNG, where gas entropy is used to quantify the effect of AGN feedback. Since high-entropy gas can have high temperature and thus high specific internal energy to be dispersed before forming stars, and have low cooling rate due to the high resistance to compression, the use of entropy as a proxy for the feedback effect is intuitive. Their results show that the entropy of the gas in the circumgalactic medium (CGM) of massive galaxies is significantly lifted by the kinetic-mode feedback out to a distance of several hundred kiloparsecs, so that quenching can be sustained. An interesting side effect of the large-scale kinetic-mode feedback, as proposed by Martín-Navarro et al. (2021), is the reduction of ram pressure stripping of satellite galaxies along the minor axes of the central galaxies where the gas is preferentially cleared by the feedback from the central galaxies. However, as clearly demonstrated by Kang et al. (2007) and Karp et al. (2023), the angular variation of subhalo properties in halos also produces significant anisotropy in the properties of satellite galaxies around central galaxies. A careful subtraction of the effects of subhalo anisotropy is thus necessary to isolate the effect of the AGN feedback on satellite galaxies at large scales.
@@ 5.1 The build-up of scaling relations
Figure 11: $_{\rm BH}-MM_{*}$ (left panel) and $M_{\rm BH}$-$M_{\rm h}$ (right panel) relations at different time points, $t_{23}$ (blue contours), $t_{\rm q}$ (red contours), and $z=0$ (black dots), in the histories of the galaxies. In each panel, three dashed lines indicate three relations with different logarithmic slopes. Contours from inner to outer enclose 10% – 95% of the sample. See §5.1 for the details about the build-up of these relations.
As discussed in §4, the evolution of a SMBH (and its host galaxy/halo) involves a series of critical times that mark the transitions under specific conditions. In the $_{\rm BH}-MM_{*}$ (or $M_{\rm BH}$-$M_{\rm h}$) plane, these conditions form a series of edges that a SMBH has to pass before it can reach subsequent phases. SMBHs with different initial conditions start at different positions in the plane, “flow” in the plane following different paths, and cross the edges at different locations and times. Thus, the $_{\rm BH}-MM_{*}$ (or $M_{\rm BH}$-$M_{\rm h}$) scaling relations at a given epoch for a population of SMBHs in a given phase can be viewed as a time slice of the flow lines in the plane between the two edges that bracket the phase in question. If observations are able to obtain such slices of different redshifts, the locations of the edges can be inferred, thus providing constraints on the physical processes that drive the transitions in the evolution.
Fig. 11 shows the distributions of SMBHs in the $_{\rm BH}-MM_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ planes at three times: $t_{23}$, when the fast accretion of SMBH ends, $t_{\rm q}$ ($\equiv t_{34}$), when the galaxy is quenched, and $z=0$. As discussed in §4.2, the transition at $t_{23}$ happens when the thermal-mode feedback of AGN is able to unbind the gas of the galaxy. According to Eqs. (14) and (4.3.1), this corresponds to a condition of
@@
which predicts a power-law relation between $M_{\rm BH}$ and $M_{*}$ (or $M_{\rm h}$), with a power index $\gamma=5/3$ at the time when the star formation efficiency peaks, or a different index if $f_{\rm gas}$ and $f_{*}$ are mass-dependent. As shown by the blue contours in the left panel of Fig. 11, the $_{\rm BH}-MM_{*}$ relation at $t_{23}$ follows a power-law with a power index of about $5/3$. This implies a nearly constant $f_{\rm gas}f_{*}^{-5/3}$ at around $t_{\rm 23}$, and suggests that the gas and stellar masses are positively correlated. At $M_{*}\approx 10^{10}\,{\rm M_{\odot}}$ and $M_{\rm BH}\approx 10^{7}\,{\rm M_{\odot}}$, there is a population of SMBHs that deviates from the relation. As noted in §3, this is mainly due to the difficulty in determining $t_{23}$ accurately. On the other hand, as shown by the blue contours in the right panel, the $M_{\rm BH}$-$M_{\rm h}$ relation in logarithmic scales appears to be non-linear even for SMBHs with $M_{\rm BH}$ well above $10^{7}\,{\rm M_{\odot}}$. This reflects the non-linear dependence of $f_{\rm gas}$ on $M_{\rm h}$. At about $M_{\rm h}=M_{\rm h,c}=10^{12}\,{\rm M_{\odot}}$, $f_{\rm gas}$ reaches its peak, leading to a $M_{\rm BH}$-$M_{\rm h}$ relation that follows the power-law of $\gamma=5/3$. At lower or higher $M_{\rm h}$, the dependence of $f_{\rm gas}$ on $M_{h}$ changes, which results in bending in the $M_{\rm BH}$-$M_{\rm h}$ relation.
@@
As concluded in §4.3, the evolution of SMBHs between $t_{23}$ and $t_{\rm q}$ is self-regulated so as to follow $M_{\rm BH}\propto M_{*}^{5/3}$. Naively one would expect that the $_{\rm BH}-MM_{*}$ relation at $t_{\rm q}$ follows the same power law. However, as shown by the red contours in the left panel of Fig. 11, the $_{\rm BH}-MM_{*}$ relation at $t\sim t_{\rm q}$ appears to be much shallower, with $\gamma<1$. As discussed in §4.3.2, the trigger of the kinetic-mode feedback is controlled by the $M_{\rm BH}$-dependent threshold of the Eddington ratio, $\chi_{\rm thres}$, for SMBHs with $M_{\rm BH}\lesssim 7\times 10^{8}\,{\rm M_{\odot}}$. It can also be seen from Fig. 8 that the majority of galaxies are quenched soon (within $\approx 1\,{\rm Gyr}$) after the trigger of the kinetic-mode feedback for halos with $M_{\rm h}<10^{12.3}\,{\rm M_{\odot}}$. In TNG, $\chi_{\rm thres}$ is set so that the kinetic-mode feedback is triggered preferentially for high-mass SMBHs. Thus, the quenching of high-mass galaxies occurs earlier than that expected from a model with a constant (and small) $\chi_{\rm thres}$. The shallow $_{\rm BH}-MM_{*}$ relation at $M_{\rm*}\lesssim 10^{11}\,{\rm M_{\odot}}$ is a direct consequence of the choice of the $\chi_{\rm thres}$-$M_{\rm BH}$ relation. For higher masses ($M_{\rm BH}\gtrsim 7\times 10^{8}\,{\rm M_{\odot}}$), the threshold $\chi_{\rm thres}$ reaches the upper limit, 0.1, and becomes independent of $M_{\rm BH}$. Meanwhile, the conditions necessary for the quenching of galaxies in high-mass halos are delayed by frequent mergers (§4.3.3). Eventually, the preference for quenching in high-mass galaxies is weakened, which allows the SMBHs in them to grow further. The tip present in the red contours at $M_{*}\gtrsim 10^{11}\,{\rm M_{\odot}}$ in the left panel of Fig. 11 verifies this. The $M_{\rm BH}$-$M_{\rm h}$ relation at $t_{\rm q}$, as shown in the right panel, appears as shallow as $\gamma=0.5$. The difference in the index between $_{\rm BH}-MM_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ relations at $t_{\rm q}$ can be understood by an argument similar to that used for the bending $M_{\rm BH}$-$M_{\rm h}$ relation at $t_{\rm 23}$, but here both $f_{\rm gas}$ and $f_{*}$ decrease with increasing $M_{\rm h}$ at $M_{\rm h}\gtrsim 10^{12}\,{\rm M_{\odot}}$.
@@
The evolution of SMBHs (and their host galaxies) in the final phase after $t_{\rm q}$ is dominated by the ex-situ growth. As shown in §4.4, the fraction of ex-situ growth for both $M_{\rm BH}$ and $M_{*}$ depends significantly on the halo mass at $z=0$. This mass dependence implies that the growth of SMBHs and galaxies in low-mass halos is almost frozen after $t_{\rm q}$, while the growth in massive halos is not. The net effect is a “stretch” of the $_{\rm BH}-MM_{*}$ relation towards the high-mass end from an anchor point at the low-mass end, as shown by the black dots in the left panel of Fig. 11. As mergers tend to give a 1:1 relation between the involved masses, the $_{\rm BH}-MM_{*}$ and $M_{\rm BH}$-$M_{\rm h}$ relations at the most massive ends steepen relative to those at $t_{\rm q}$. The history of the scaling relations is then erased by mergers, leaving behind little trace of the seeding and early growth of the SMBHs.
@@
The TNG subgrid model for AGN feedback is designed to reproduce the observed $M_{\rm BH}$-$M_{\rm*,bulge}$ relation of Kormendy & Ho (2013), which is a power-law form with an index of $\gamma\approx 1$. As shown by Weinberger et al. (2017, see their figure 5), the $M_{\rm BH}$-$M_{\rm*,bulge}$ relation of the TNG model roughly follows the $\gamma=1$ relation, and so does the $_{\rm BH}-MM_{*}$ relation. Since current observational results still have significant uncertainties in estimating the SMBH and bulge masses, it is still unclear whether or not the observed $_{\rm BH}-MM_{*}$ (or $M_{\rm*,bulge}$) relation can be extended to lower SMBH masses, and what the power index really is. In a more recent work, Greene et al. (2020) suggested a steeper relation, $M_{\rm BH}\sim M_{*}^{1.33}$ for early-type galaxies. Graham & Sahu (2023a, see their table 2) obtained $M_{\rm BH}\sim M_{*}^{1.69}$ (and $M_{\rm BH}\sim M_{\rm*,bulge}^{1.64}$) for pure elliptical galaxies. For the $M_{\rm BH}$-$M_{\rm h}$ relation, Voit et al. (2024) showed that the prediction from TNG is significantly flatter than observations. For the model to match the observations, they suggested modifications of the criteria for the triggering of the kinetic-mode feedback and/or the feedback efficiency. These more recent observations in the nearby Universe can thus put constraints on the subgrid physics to be implemented in the next generation of simulations, especially for the SMBH growth in the phase of quenched evolution.
@@
Since mergers in the quenched phase erased the imprints of early evolution, in particular for SMBHs in massive halos, a slice of the $_{\rm BH}-MM_{*}$ (or $M_{\rm BH}$-$M_{\rm h}$) flow lines in the nearby Universe may not be adequate to infer the full tracks of galaxies. This can be seen by comparing our TNG-based conclusions to those inferred by Graham & Sahu (2023a, see their figures 8 and 9) from the observed $M_{\rm BH}$-$M_{\rm*,bulge}$ relations for galaxies with different morphological types. At $M_{\rm*,bulge}\gtrsim 10^{11}\,{\rm M_{\odot}}$, their analysis suggests that mergers between elliptical galaxies (with bulge-to-total ratio $\approx 1$) tend to drive migrations in the $M_{\rm BH}$-$M_{\rm*,bulge}$ plane towards a 1:1 relation, consistent with our inferences from the TNG. However, our analysis suggests that, for galaxies with $M_{*}\lesssim 10^{10.5}\,{\rm M_{\odot}}$, the ex-situ growth in both $M_{*}$ and $M_{\rm BH}$ contributes little to the total mass, which is inconsistent with the merger-driven scenario for the formation of elliptical galaxies and the build-up of the $M_{\rm BH}$-$M_{\rm*,bulge}$ relation in the low-mass end. One possibility is that the dynamical hotness of these galaxies comes from the fast inflow of gas associated with the fast regime of halo assembly (e.g. Zhao et al., 2003), as suggested by Mo et al. (2024, see also ). The change of the power index of the $_{\rm BH}-MM_{*}$ relation at $t_{\rm q}$, as demonstrated in this paper, is then expected to occur if the kinetic-mode feedback drives the quenching of galaxies.
@@
On the other hand, observations directly targeting SMBHs at high $z$, where they are less massive and more active, are crucial to cover the early transitions/edges in the evolutionary history of the SMBH growth and to avoid the ambiguity in the extrapolation of the local scaling relations. Using the deep imaging and spectroscopy capabilities of JWST, Wang et al. (2024b) found a galaxy with a spectrum compatible with that produced by a Pop-III star cluster. As theoretical studies (e.g. Stacy et al., 2016; Wise et al., 2019; Chon & Omukai, 2020; Latif et al., 2022a, b; Regan & Volonteri, 2024; Ventura et al., 2024) have suggested that massive Pop-III stars are promising SMBH seeds of intermediate masses ($\sim 10^{4}\,{\rm M_{\odot}}$), direct observations of Pop-III stars (clusters) will be crucial to distinguish among different seeding scenarios (see, e.g. Spinoso et al., 2023) and to probe the growth of SMBHs in Phase-1.